OpenStudy (anonymous):

Mrs. Jones is 28 years older than her daughter. In 11 years she will be 3 times as old as her daughter. How old is each now?

OpenStudy (anonymous):

Let Mrs Jones be represented as J. and the daughter be represented as D.\[J = D + 28\]\[J+11=(D+11)\times3\] Do you understand how I made these equations?

OpenStudy (anonymous):

Yes, I do understand

OpenStudy (anonymous):

Great solve for J and that's your answer. You could start by simplifying the second equation